The traffic light turns green and the driver of a high-performance car slams the accelerator to the floor. The accelerator registers 22 m/s2. Convert this reading to km/hr2
1
Expert's answer
2020-06-10T18:25:28-0400

The conversation will be the following:


"a = 22\dfrac{m}{s^2} =22\cdot \dfrac{ 10^{-3}km}{(1/3600)^2hr^2} = 22\cdot12960\dfrac{km}{hr^2} = 285120\dfrac{km}{hr^2}"

Answer. a = 285120 km/hr^2